Solve coshz=2 and separate real and Imaginary part?
1
Expert's answer
2020-10-17T14:50:12-0400

"cosh z=2\Rightarrow e^z+e^{-z}=4" . Let "e^z=a." Then "a+\frac{1}{a}=4\Rightarrow a^2-4a+1=0." Hence "a= 2\pm\sqrt{3} ." Hence "e^z= 2\pm\sqrt{3}." Now if "z=x+iy\Rightarrow e^x(cos y+i sin y)=2\pm\sqrt{3}." As it has no imaginary part, so "y=n\pi." Since RHS ">0," "e^xcosy >0" . Hence "n=2m." "\therefore e^x=2\pm\sqrt{3}\Rightarrow x=ln(2\pm\sqrt{3})" "\therefore" "z=ln(2\pm \sqrt{3})+i2m\pi."